OpenStudy (usukidoll):

Define a product of pairs setting \((a,b) \cdot (c,d) = (ac+bd,ad+bc)\) . Show that this operation is well defined for equivalence classes.

OpenStudy (usukidoll):

My attempt: By definition 6.2.9, let \(R\) be an equivalence relation on a set \(S\). For each element \(x \in S\) \([x]=[y \in S: (x,y) \in R]\) is the equivalence relation and partition. Suppose \((a,b)\) or (a',b')\( and \)(c,d)\( or \)(c',d')\(. Then we need to show that \)((ac+bd+a'd'+b'c'=a'c'+b'd'+ad+bc))\( \)ac+bd+a'd'+b'c'=a'c'+b'd'+ad+bc\( Adding \)bc'+a'c'\( to both sides, we have \)ac+bd+a'd'+b'c'=a'c'+b'd'+ad+bc\(.

OpenStudy (usukidoll):

so what I don't get is where does this \[(a,b) ~ (a'b') \rightarrow a+b'=b+a'\] come into play... I feel like the reflexivity definition is involved

OpenStudy (anonymous):

Latex spaghetti

OpenStudy (usukidoll):

hold on

OpenStudy (usukidoll):

My attempt: By definition 6.2.9, let \(R\) be an equivalence relation on a set \(S\). For each element \(x \in S\) \([x]=[y \in S: (x,y) \in R]\) is the equivalence relation and partition. Suppose \((a,b)\) or \((a',b')\) and \((c,d)\) or \((c',d')\). Then we need to show that \(((ac+bd+a'd'+b'c'=a'c'+b'd'+ad+bc))\) \(ac+bd+a'd'+b'c'=a'c'+b'd'+ad+bc\) Adding \(bc'+a'c'\) to both sides, we have \(ac+bd+a'd'+b'c'=a'c'+b'd'+ad+bc\).

OpenStudy (usukidoll):

but before the adding part suppose ... there's this part.. \[(a+b')c' = (b+a')c'\] which becomes \[ac'+b'c'=bc'+a'c'\] so where does that come from because apparently I need to add that to both sides and cancel out b'c' and a'c' from\[ac+bd+a'd'+b'c'+bc'+a'c'=a'c'+b'd'+ad+bc+ac'+b'c'\]

OpenStudy (usukidoll):

so after cancelling... we have \[ac+bd+a'd'+bc'=b'd'+ad+bc+ac'\] From \[(a+b')d'=(b+a)d'\] we have\[ad'+b'd'=bd'+a'd'\] so adding bd'+a'd' to both sides yields, \[ac+bd+a'd'+bc'+ad'+b'd'=b'd'+ad+bc+ac'+bd'+a'd'\]

OpenStudy (anonymous):

That definition is confusing

OpenStudy (anonymous):

Can you just restate the definition again?

OpenStudy (usukidoll):

cancelling, a'd' and b'd' we have \[ac+ad'+bc'+bd=ad+bc+ac'+bd'\] factoring out a and b , we have \[a(c+d')+b(c'+d) = a(d+c')+b(c+d')\] there is an equivalence relation on the left and the operation is well defined on the right.

OpenStudy (usukidoll):

By definition 6.2.9, let $R$ be an equivalence relation on a set $S$. For each element $x \in S$ $[x]=[y \in S: (x,y) \in R]$ is the equivalence relation and partition.

OpenStudy (usukidoll):

grtrr

OpenStudy (usukidoll):

By definition 6.2.9, let \(R\) be an equivalence relation on a set \(S\). For each element \(x \in S\) \([x]=[y \in S: (x,y) \in R]\)

OpenStudy (anonymous):

So \([x]\) is an equivalence class, right?

OpenStudy (anonymous):

This is the definition of an equivalence class?

OpenStudy (usukidoll):

all that letter gibberish stuff is what my prof wrote... and that stuff is correct... but what I don't get is how the heck does this happen \[(a,b) ~ (a',b') \rightarrow a+b'=b+a'\] (a,b) ~(a'b')

OpenStudy (usukidoll):

feels like reflexivity

OpenStudy (anonymous):

I cannot help, because I don't understand the problem.

OpenStudy (anonymous):

What does it mean for an operation to be well defined for equivalence classes?

OpenStudy (usukidoll):

same goes for \[(a+b')c'=(b+a')c' \rightarrow ac+b'c'=bc'+a'c'\] \[(a+b')d'=(b+a')d' \rightarrow ad+b'd'=bd'+a'd'\] where dd that come from ?!

OpenStudy (usukidoll):

an expression is well-defined if it is unambiguous and its objects are independent of their representative. :/

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That is not a very formal definition

OpenStudy (anonymous):

How would you prove something is unambiguous?

OpenStudy (usukidoll):

???

OpenStudy (anonymous):

You gave a definition for products of pairs...

OpenStudy (anonymous):

Yet you don't have any pairs anywhere....

OpenStudy (anonymous):

I still don't know the formal definition, of even any examples and counter examples, of what a well defined operation are.

OpenStudy (anonymous):

Maybe they mean it is closed?

OpenStudy (anonymous):

\((a,b)\in R\) and \((c,d)\in R\) We do: \[ (a,b)\cdot (c,d)=(ac+bd,ad+bc) \]Is \( (ac+bd,ad+bc) \in R\)?

OpenStudy (anonymous):

I guess we have to explore: reflexivity symmetry transitivity

OpenStudy (zzr0ck3r):

well defined \(a=b\implies f(a) = f(b)\)

OpenStudy (anonymous):

Hmmmmm, well defined applies to operations? What would be not well defined?

OpenStudy (zzr0ck3r):

operations are functions

OpenStudy (zzr0ck3r):

XxX -> X

OpenStudy (anonymous):

But functions seems to be well defined by their definition as is based on your definition of well defined.

OpenStudy (zzr0ck3r):

definition of function includes well defined and defined everywhere

OpenStudy (zzr0ck3r):

\(f:\mathbb{Q}\rightarrow R, f(\frac{a}{b})=b\) is not well defined relation

OpenStudy (zzr0ck3r):

\(\frac{2}{4}=\frac{1}{2}\) but \(4\ne2\)

OpenStudy (zzr0ck3r):

anyway i didn't read what you guys where doing, I just thought you wanted exact definition of well defined. The only time you even have a problem with something being or not being well defined is when dealing with equivalence classes.

OpenStudy (anonymous):

I think your definition helps quite a bit.

OpenStudy (zzr0ck3r):

so you need to show that if \((a,b) = (c,d)\) and \((a',b')=(c',d')\) then \((a,b)\cdot(a',b')=(c,d)\cdot(c',d')\)

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

\[ (a,b)\cdot (a',b') = (aa'+bb',ab'+ba') \]

OpenStudy (zzr0ck3r):

I just had a similar topic in class today trying to show that coset multiplication is well defined operation on the set of equivalence classes. its fresh:)

OpenStudy (anonymous):

\[ (c,d)\cdot (c',d') =(cc'+dd',cd'+dc') \]

OpenStudy (anonymous):

We need to show that \(aa'+bb' = cc'+dd'\) then? And also that \(ab'+ba' = cd'+dc'\)?

OpenStudy (zzr0ck3r):

yeah, using the fact that (a,b) = (c,d), and (a',b')=(c',d') and what that is defined to mean

OpenStudy (zzr0ck3r):

im on my phone so some of the code is showing weird where he defined things. Ill go get the wife's pc.

OpenStudy (zzr0ck3r):

we still need to figure this out?